3. A turbine has a power output of 7500 kJ/sec when steam with a mass flow rate of 11.11
kg/sec enters the turbine at 773 K, 8MPa and leaves as saturated vapor at 40 kPa. If the
surrounding of the turbine is at 25 °C, find the rate of entropy generation at steady state.
1
Expert's answer
2020-11-23T10:32:34-0500

The steam mass flow rate m = 11.11 kg/sec

P1 = 8 MPa

T1 = 773 K

P2 = 40 kPa

"W_{out} = 7500 \;kJ/sec"

"T_{sur} = 25 + 273 = 298 \;K"

h1 = 3399.5 kJ/kg

s1 = 6.7266 kJ/kgK

And at 40 kPa

h2 = hg = 2636.1 kJ/kg

s2 = sg = 7.6691 kJ/kgK

"m_{in} - m_{out} = \u2206m = 0"

"m_1 = m_2 = m"

"E_{in} - E_{out} = \u2206E = 0"

"E_{in} = E_{out}"

"mh_1 = Q_{out} + W_{out} + mh_2"

"Q_{out} = m(h_1 - h_2) - W_{out}"

"Q_{out} = 11.11(3399.5 - 2636.1) - 7500 = 981.4 \;kJ/s"

"S_{in} - S_{out} + S_{gen} = \u2206S_{sys} = 0"

"ms_1 - ms_2 - \frac{Q_{out}}{T_{sur}} + S_{gen} = 0"

"S_{gen} = m(s_2 - s_1) + \frac{Q_{out}}{T_{sur}}"

"S_{gen} = 11.11(7.6691 - 6.7266) + \frac{981.4}{298} = 10.47 + 3.29 = 13.76 \;kW/K"